# Constitutional Consensus

### v4

@@ 5.4 Epoch Transitions via Constitutional Amendments
Amendment Finalization by $p$: Upon issuing an ordinary block, if $p$ knows of a valid constitutional amendment decision $d$ that is not final in $B$ then issue the block with $d$ instead of with payload. End Epoch by $p$: Upon finalising in $B$ a constitutional block $b$ with constitutional amendment $d$, prevailing population $P$, and new population $P^{\prime}$, call $\tau(b)$, create a $(\textsc{coronate},d)$ block, send it to members of $P\cup P^{\prime}$, and cease participation in the epoch. Start Epoch by $q$: Let $d$ be a valid decision to amend the prevailing constitution, with new population $P^{\prime}$.
@@
Respond to Belated nack by $q$: Upon receipt of a nack $q^{\prime}$-block for a $q$-block $b$ of a past epoch $e$ for which $q^{\prime}\in\Pi$ is a participant, judiciously send $[b]$ to $q^{\prime}$.
@@
In the good case, a new constitutional amendment $d$ is issued by the leader of a wave and finalized within that same wave. In the worst case, it might take arbitrarily many waves to finalize $d$ prior to GST or due to faulty leaders. After GST, a constitutional amendment block issued by a correct leader among $P$ is bound to be finalized (and such a leader is bound to be selected eventually due to the fairness of $\textit{leader}_{P}$). Once finalized, correct members of $P$ will issue a coronate block and correct members of $P^{\prime}$ will receive it, allowing the current epoch to end and the next one to begin. This argument informally establishes the liveness of the transition among constitutions due to constitutional amendments.
5.5 Safety and Liveness of Constitutional Amendments
Theorem 5.4 (Constitutional Amendment Correctness).
@@ 5.5 Safety and Liveness of Constitutional Amendments
We need to prove that in every run of the protocol, there exists a sequence $T$ s.t. the two safety properties and three liveness properties of Definition 2.4 hold w.r.t. $T$. Note that the $i^{th}$ epoch $e_{i}$ is in fact $d_{i}$. We have proven that Algorithm 1 is correct in each epoch. Therefore, there is a sequence $T^{i}$ so that the output sequences produced by all participants in epoch $e_{i}$ are consistent and valid w.r.t. $T^{i}$.
@@
Lemma 5.5 (Constitutional Amendment Safety).
@@
Proof 5.6.
Validity immediately follows from the fact that agents participate only in epochs of which they are members. Consistency follows from the fact that agents participate in epochs sequentially, namely, they stop running the protocol in epoch $i$ before starting epoch $j$.
Lemma 5.7 (Constitutional Amendment Liveness).
@@
Proof 5.8.
Consider an epoch $e$ with prevailing constitution $(P,\sigma,\Delta)$, and the two liveness requirements. If $e$ is the last epoch then Epoch Liveness follows from the liveness of the single-epoch Algorithm 1, Theorem 4.10. Else assume $e$ is not the last epoch, and therefore it must have ended with a coronation by a $\sigma$-supermajority among $P$, all observing a wave finalizing a constitutional block $b$ with payload $d$.
Let $p\in P$ be a correct agent. We argue that $p$ must eventually coronate $d$, and while doing so, call $\tau(b)$. As all correct participants in the epoch $e$ end it by calling $\tau(b)$, all would output the same blocks for $e$. If $p$ is among the $\sigma$-supermajority coronating $d$ we are done. Else there is at least one correct agent $q\in P$ that coronated $d$ and sent to $p$ a coronate block $b^{\prime}$ observing a wave $W$ that finalizes $b$ . Since $q$ and $p$ are correct, eventually a $\textsc{nack}(b^{\prime})$-block by $q$ will be received by $p$, who will respond (even after the epoch $e$ has ended, thanks to the Respond to Belated nack rule) by sending to $q$ any blocks in $W$ and their predecessors that $q$ is missing, upon receipt of which $p$ would issue a $(\textsc{coronate},d)$ block and output $\tau(b)$.
Considering Constitution Amendment Liveness, once a constitutional amendment $d$ is known, all correct participants follow the high-throughput protocol and produce only blocks with payload $d$. Eventually a correct leader will finalize its block, following the same argument as in the proof of Theorem 4.10, upon which all correct participants will eventually issue $(\textsc{coronate},d)$ blocks.
6 Constitutional Consensus is Grassroots
We recall the formal definition of a grassroots protocol from reference [38], instantiate it informally (bypassing the formalities of the definition in [38]) to the Constitutional Consensus protocol, and then argue that the protocol is grassroots.
Definition 6.1 (Oblivious, Interactive, Grassroots).
@@ 6 Constitutional Consensus is Grassroots
Note that the two requirements cannot even be expressed without in the context of the basic (no constitutional amendments) Constitutional Consensus protocol, as the population $P$ is given and fixed. Also note that Constitutional Consensus being interactive critically-depends on the ability of the prevailing population $P$ to decide to add new members.
7 Conclusions
The Constitutional Consensus protocol and its extensions are novel in several respects, all devised in order to serve the needs of future grassroots platforms that operate digital social contracts. Initially, we expect low-throughput applications to prevail, as small digital communities engaged in constitutional democratic conduct use the platform to grow and evolve [39]. Even small community banks and cooperatives [37] may still be, at least initially, on the low-throughput side. As communities federate to form ever-larger communities, and local community banks federate to form regional and national banks, higher-throughput execution would follow. Until then, simulating the various scenarios would be productive future work.

### v5

@@ Appendix G Constitutional Amendment Pseudocode
Amendment Finalisation by $p$:: Upon issuing an ordinary block, if $p$’s amendment protocol has produced a valid constitutional amendment decision $d$ that is not final in $B$ then issue the block with $d$ instead of with payload. End Epoch by $p$:: Upon finalising in $B$ a constitutional block $b$ with constitutional amendment $d$, prevailing population $P$, and new population $P^{\prime}$, call $\tau(b)$, create a $(\textsc{coronate},d)$ block, send it to $P\cup P^{\prime}$, and cease participation in the epoch. Start Epoch by $q$:: Let $d$ be a valid decision to amend the prevailing constitution, with new population $P^{\prime}$.
@@
Respond to Belated nack by $q$:: Upon receipt of a nack $q^{\prime}$-block for a $q$-block $b$ of a past epoch $e$ for which $q^{\prime}\in\Pi$ is a participant, judiciously send $[b]$ to $q^{\prime}$.
@@
Epoch transition walkthrough. Algorithm 2 shows the additional logic supporting amendments in Algorithm 1. Every epoch commences with a constitutional genesis block that specifies its prevailing constitution.
Amendement finalisation: If, during the epoch, members of the prevailing population produce a valid decision $d$ to amend the prevailing constitution, they endeavour to finalise this decision by producing only constitutional blocks with $d$ as payload.
End Epoch: Once an agent finalises a constitutional block with $d$, it produces a coronate block with payload $(\textsc{coronate},d)$, sends it to both the old and new populations of $d$, and ceases participation in the epoch. Members of the old population that receive it and have yet to finalise $d$ can nack any missing dependent blocks. After ceasing to participate in the old epoch, an agent no longer adds blocks from that epoch to its blocklace. However, it must continue to respond to Belated nack requests from agents who are missing blocks in order to finalise the old epoch (due to Byzantine equivocation). An old epoch may be garbage-collected upon learning that all agents have coronated an ensuing epoch.
Start Epoch: A member of the new population that receives a $\sigma$-supermajority of coronate blocks may commence the new epoch with $d$ as the constitutional genesis block. An agent that is a member of both the new and old populations commences the new epoch after both receiving a $\sigma$-supermajority of coronate blocks and sending one of its own.
The algorithm’s output includes constitutional blocks with constitutional amendment decisions as payload. When $\tau$ orders a constitutional block with amendment $d$ for the first time, the protocol outputs the new constitution, suppressing subsequent copies.
Appendix H Safety and Liveness of Constitutional Amendments
Theorem H.1 (Constitutional Amendment Correctness).
@@ Appendix H Safety and Liveness of Constitutional Amendments
We need to prove that in every run of the protocol, there exists a sequence $T$ s.t. the two safety properties and two liveness properties of Definition 2.3 hold w.r.t. $T$. Note that the $i^{th}$ epoch $e_{i}$ is in fact $d_{i}$. We have proven that Algorithm 1 is correct in each epoch. Therefore, there is a sequence $T^{i}$ so that the output sequences produced by all participants in epoch $e_{i}$ are consistent and valid w.r.t. $T^{i}$.
@@
Lemma H.2 (Constitutional Amendment Safety).
@@
Proof.
Validity immediately follows from the fact that agents participate only in epochs of which they are members. Consistency follows from the fact that agents participate in epochs sequentially, namely, they stop running the protocol in epoch $i$ before starting epoch $j$. ∎
Lemma H.3 (Constitutional Amendment Liveness).
@@
Proof.
Consider an epoch $e$ with prevailing constitution $(P,\sigma,\Delta)$, and the two liveness requirements. If $e$ is the last epoch then Epoch Liveness follows from the liveness of the single-epoch Algorithm 1, Theorem E.10. Else assume $e$ is not the last epoch, and therefore it must have ended with a coronation by a $\sigma$-supermajority among $P$, all observing a wave finalizing a constitutional block $b$ with payload $d$.
Let $p\in P$ be a correct agent. We argue that $p$ must eventually coronate $d$, and while doing so, call $\tau(b)$. As all correct participants in the epoch $e$ end it by calling $\tau(b)$, all would output the same blocks for $e$. If $p$ is among the $\sigma$-supermajority coronating $d$ we are done.
Else there is at least one correct agent $q\in P$ that coronated $d$ and sent to $p$ a coronate block $b^{\prime}$ observing a wave $W$ that finalizes $b$. Since the network is reliable, $p$ receives $b^{\prime}$ but may be missing some blocks in $b^{\prime}]$. Then $p$ sends a $\textsc{nack}(b^{\prime})$-block to $q$, who responds (even after epoch $e$ has ended, thanks to the Respond to Belated nack rule) by sending to $p$ the missing blocks in $W$ and their predecessors, upon receipt of which $p$ finalizes $b$, issues a $(\textsc{coronate},d)$ block, and outputs $\tau(b)$.
Considering Constitution Amendment Liveness, once a constitutional amendment $d$ is known, all correct participants follow the high-throughput protocol and produce only blocks with payload $d$. Eventually a correct leader will finalize its block, following the same argument as in the proof of Theorem E.10, upon which all correct participants will eventually issue $(\textsc{coronate},d)$ blocks. ∎
Appendix I Constitutional Consensus is Grassroots
We recall the formal definition of a grassroots protocol from reference (Shapiro, 2026), instantiate it informally (bypassing the formalities of the definition in (Shapiro, 2026)) to the Constitutional Consensus protocol, and then argue that the protocol is grassroots.
Definition I.1 (Oblivious, Interactive, Grassroots).
@@ Appendix I Constitutional Consensus is Grassroots
Note that the two requirements cannot even be expressed in the context of the basic (no constitutional amendments) Constitutional Consensus protocol, as the population $P$ is given and fixed. Also note that Constitutional Consensus being interactive critically-depends on the ability of the prevailing population $P$ to decide to add new members.
Appendix J Related Work
Constitutional Consensus combines three properties: (i) democratic one-person-one-vote governance, (ii) decentralised BFT consensus, and (iii) self-governing constitutional amendment within the protocol. We survey prior work across five areas, corresponding to the five aspects of the protocol described in the introduction, and find that every existing system lacks at least one of these properties.
J.1. Smart Contracts and Digital Social Contracts
Deployed smart contract bytecode on Ethereum is immutable; amendments are achieved through proxy schemes that separate mutable state from swappable logic, controlled by admin keys, multisigs, or token-weighted governance (De Filippi et al., 2021). Tezos (Goodman, 2014) embeds self-amendment at the protocol level, and the amendment mechanism is itself subject to amendment. In both cases, governance is plutocratic and amendments apply to code, not to a broader protocol constitution.
A digital social contract (Cardelli et al., 2020) differs architecturally: it is a voluntary agreement among people, executed on the participants’ own devices, and enforced by the participants themselves. Constitutional Consensus is designed for such contracts. Governance is democratic (one person, one vote) rather than plutocratic (one token, one vote), and amendments apply to the protocol constitution—including the participant set, supermajority threshold, and timing—not merely to code. These distinctions are summarised in Table 1.
Table 1. Comparing Smart Contracts and Digital Social Contracts.
Smart Contract (De Filippi et al., 2021) Digital Social Contract (Cardelli et al., 2020)
Among: Accounts People
Executed by: Third-party validators The people’s smartphones
Cost: Gas None
Consensus: Always required Only when required by the contract
Governance: Plutocratic (token-weighted voting) Democratic (one person, one vote)
Amended via: Proxy schemes and redeployment; bytecode is immutable Constitutional amendments
Amendment authority: Token-weighted governance Participants themselves via supermajority
Applications: Decentralized Finance (DeFi) (Schueffel, 2021; Jensen et al., 2021),
Decentralized Autonomous Organizations (DAOs) (Ethereum, ndao; Ding et al., 2023),
Non-Fungible Tokens (NFTs) (Wang et al., 2021)
Social graph and social networking (Shapiro, 2023b),
grassroots currencies (Shapiro, 2024; Lewis-Pye et al., 2023),
democratic communities and their federation (Shapiro and Talmon, 2025)
J.2. DAOs and Blockchain Governance
Decentralised Autonomous Organisations (Hassan and De Filippi, 2021) are the prevailing method for participant-governed digital organisations. Several blockchains can modify their own protocol rules through on-chain stakeholder voting: Tezos (Goodman, 2014) through a five-period procedure requiring an 80% supermajority; Polkadot (Wood, 2016) through forkless runtime upgrades via on-chain referenda; Cardano (Corduan et al., 2023) through a tripartite structure of Constitutional Committee, Delegated Representatives, and Stake Pool Operators; Cosmos (Kwon and Buchman, 2016) through its governance module; and the Internet Computer (Camenisch et al., 2022) through its Network Nervous System. All are universally plutocratic: [^5] Cardano is utilizing one-person-one-vote but only for its Constitutional Committee, which has essentially only veto power. voting power is proportional to staked tokens. This is not accidental; stake-weighting provides built-in Sybil resistance, whereas democratic voting requires an external mechanism to establish unique personhood.
DAO frameworks built on existing blockchains are similarly plutocratic. Aragon (Cuende and Izquierdo, 2017) defaults to ERC-20 token-weighted voting. Compound Governor (Leshner, 2020) and OpenZeppelin Governor (OpenZeppelin, 2021) are built around ERC20Votes, tying governance power to token holdings. Snapshot (Marino and Snapshot Labs, 2020) offers configurable voting strategies including an equal-weight option, but is purely off-chain and admin-controlled. Optimism’s Citizens’ House (Optimism Foundation, 2022) is a notable exception: it implements one-person-one-vote governance for retroactive public goods funding. However, the Citizens’ House relies on Ethereum’s proof-of-stake consensus, and the Optimism Foundation retains significant authority over proposal execution.
Schneider (Schneider, 2022) argues that the economic logic embedded in blockchain systems systematically constrains democratic possibilities. Siddarth et al. (Siddarth et al., 2020) identify the gap directly: “the lack of a robust notion of personhood … has led to the development of plutocracies.” Talmon (Talmon, 2023) examines DAO governance from a computational social choice perspective. Buterin (Buterin, 2021) critiques coin voting as creating plutocratic governance vulnerable to vote buying and proposes proof-of-personhood alternatives. No DAO or self-amending blockchain achieves democratic governance with its own BFT consensus and self-governing constitutional amendment.
Consortium blockchain platforms such as Hyperledger Fabric (Androulaki et al., 2018) and R3 Corda (Brown et al., 2016) implement organisational governance rather than democratic governance. In Fabric, configuration changes—including membership, block parameters, and governance policies—require signatures from a majority of organisation administrators, collected off-chain and submitted as configuration update transactions. Corda centralises governance further: a network operator controls certificate issuance and network parameters. While Fabric’s hierarchical policy system bears structural resemblance to constitutional amendment (each policy specifies who can modify it), decision-making authority rests with designated administrators, not with individual participants exercising equal voting rights.
J.3. Reconfiguration of Consensus Protocols
Constitutional amendments generalise reconfiguration in two respects:
  1. (1)
    Scope. Reconfiguration refers to changing the set of active agents (servers, processes), whereas a constitution describes the entire consensus protocol, and in principle any aspect of it may be amended.
  2. (2)
    Authority. Even in self-governing systems like Raft (Ongaro and Ousterhout, 2014), reconfiguration is a mechanism available to whoever holds the leader role. Constitutional amendments, by contrast, are decided by the participants themselves through a process specified by the constitution.
In Constitutional Consensus, constitutional amendment preserves the protocol instance: the transaction history (the blocklace) carries over from one epoch to the next under the new constitution. Starting a new instance, by contrast, would discard all prior state. In this work, the amendable component of the constitution consists of $(P,\sigma,\Delta)$ (Definition 2.1), parameters which affect the safety, liveness, and performance of almost any consensus protocol for eventual synchrony.
[…]
The reconfiguration literature has focused on how membership changes execute safely. Vertical Paxos (Lamport et al., 2009) introduced an auxiliary configuration master that determines the acceptor set for each configuration. DynaStore (Aguilera et al., 2011) allows any client to propose reconfigurations but treats them as externally triggered. Spiegelman et al. (Spiegelman et al., 2017) explicitly model an administrator invoking reconfiguration operations. SMART (Lorch et al., 2006) handles crash-fault reconfiguration via an external coordinator. Raft (Ongaro and Ousterhout, 2014) treats configuration changes as special log entries committed through joint consensus of both old and new configurations. In the BFT setting, BFT-SMaRt (Bessani et al., 2014) provides the first practical BFT library with reconfiguration support, but relies on an external trusted third party for membership decisions. Duan and Zhang (Duan and Zhang, 2022) showed that prior BFT reconfiguration approaches can compromise liveness, and introduced Dyno, the first formally analysed dynamic BFT protocol. Proof-of-Stake protocols achieve a form of self-reconfiguration, where validator sets change based on staking transactions (Komatovic et al., 2025), but governance is plutocratic.
ComChain (Vizier and Gramoli, 2020) and the subsequent SMaRtChain (Bessani et al., 2020) achieve BFT reconfiguration without a trusted third party: current participants run Byzantine consensus to decide on membership changes and employ a forgetting protocol that decouples permanent keys from per-view consensus keys to prevent removed nodes from creating forks. However, neither system provides a governance mechanism: the criteria for accepting new members are application-specific (e.g., proof-of-work, certification by an authority). Moreover, both are limited to membership changes and do not support amending protocol parameters or governance rules.
Across all of these, decision authority is either left unspecified, delegated to an external entity, determined by application-specific criteria, or determined plutocratically. The mechanism of reconfiguration is well-studied; the question of who decides—through democratic governance—is not.
J.4. BFT Consensus Protocols
PBFT (Castro and Liskov, 1999) introduced the first practical algorithm for BFT state machine replication, tolerating up to $\lfloor(n-1)/3\rfloor$ Byzantine faults. HotStuff (Yin et al., 2019) achieved linear communication complexity with responsiveness and became the basis for the Diem blockchain. Tendermint (Buchman, 2016) brought BFT consensus to blockchains and underlies the Cosmos network.
A second line of work replaces the leader-based approach with DAG-based protocols. DAG-Rider (Keidar et al., 2021) is the first asynchronous BFT Atomic Broadcast protocol achieving optimal resilience, amortised communication complexity, and time complexity simultaneously. Narwhal and Tusk (Danezis et al., 2021) separate reliable transaction dissemination from ordering, enabling high throughput. Bullshark (Giridharan et al., 2022) adds a practical low-latency synchronous fast path. Jolteon and Ditto (Gelashvili et al., 2022) combine leader-based consensus with an asynchronous DAG fallback. Mysticeti (Babel et al., 2023) achieves the theoretical lower bound of three message delays for commit latency. Raikwar et al. (Raikwar et al., 2024) provide a systematisation of knowledge on DAG-based consensus.
Constitutional Consensus builds on two protocols from this line of work: Cordial Miners (Keidar et al., 2023), a DAG-based BFT Atomic Broadcast protocol, and Morpheus (Lewis-Pye and Shapiro, 2025), which adapts between quiescent and high-throughput modes. Both are used as building blocks; Constitutional Consensus adds constitutional amendment and democratic governance above them. None of the BFT protocols surveyed here addresses governance or self-amendment: they provide the consensus mechanism but leave the question of who participates and how rules change to external processes.
J.5. Grassroots Protocols and Democratic Identity
Democratic digital governance platforms achieve one-person-one-vote participation but rely on centralised infrastructure. Decidim (Barandiaran et al., 2024), used by over 500 institutions, runs as a client-server web application. Loomio (Loomio Cooperative Limited, 2024) provides equal-weight deliberation and voting, but all data resides on centralised servers. CONSUL Democracy (Royo et al., 2020; Consul Democracy Foundation, 2025), serving approximately 250 governments, authenticates users via national ID but operates as a centralised web application. LiquidFeedback (Behrens et al., 2018) published a blockchain prototype with a swarm-behaviour-inspired consensus algorithm where node influence derives from the number of accredited persons trusting each node, and the roadmap envisions voting on application code. However, LiquidFeedback implements liquid democracy with transitive delegation rather than one-person-one-vote; its consensus is not characterised as BFT; and the self-amendment capability remained a roadmap item.
Proof-of-personhood (PoP) systems solve the Sybil resistance prerequisite for democratic governance. Ford (Ford, 2020) argues that digital personhood—inalienable participation rights built on proof-of-personhood mechanisms—is a necessary foundation for digital democracy. PoPCoin (Borge et al., 2017) uses pseudonym parties to enable equal-weight participation in ByzCoin consensus, achieving democratic identity with BFT consensus but not addressing self-governing amendment. Humanode (Kavazi et al., 2023) is the closest existing system to combining all three properties: it runs BFT consensus (BABE/GRANDPA on Substrate) with equal per-node weight, using biometric identity verification. However, its self-governance mechanism (Vortex DAO) operates as an off-chain simulation as of early 2026, with full authority transfer planned approximately four years after mainnet launch; Vortex is designed as a DAO overlay on the consensus protocol, not as constitutional amendment embedded within it. Idena (Idena Team, 2018) achieves equal-weight consensus via flip-test identity verification, but its consensus is not formally characterised as BFT, and protocol upgrades are controlled by the development team. Encointer (Brenzikofer, 2019) provides proof of personhood through physical key-signing ceremonies but operates as a Polkadot parachain, inheriting its consensus.
Because Constitutional Consensus requires no external authority, global ledger, or staked capital, multiple independent instances can form, operate, and grow concurrently—the grassroots property (Shapiro, 2023a, 2026). This contrasts with Proof-of-Stake protocols like Ethereum 2.0 (Buterin et al., 2020), whose security depends on the staked token having real-world value, creating a winner-take-all dynamic that precludes independent instances. Democratic platforms achieve one-person-one-vote but are centralised; PoP systems solve identity but remain building blocks. To the best of our knowledge, no existing system combines democratic one-person-one-vote governance, decentralised BFT consensus, and self-governing constitutional amendment into a single protocol.